# Adaptive Multilevel Splitting: First Application to Rare-Event Derivative Pricing

### v1

@@ 2.3.1 Theoretical properties of AMS
Well-posedness and termination.
Let $X=(X_{t})_{t\geq 0}$ be a Markov process with importance function $\xi$ and rare set $D$. For fixed $n$ and $k\in\{1,\dots,n-1\}$, AMS is well-posed: the cutting level $Z$ is an order statistic and, under standard assumptions (Feller property of $X$, continuity of $\xi$, strict entrance condition), the algorithm terminates almost surely after finitely many iterations [24].
@@
The estimator
$\hat{p}_{\mathrm{AMS}}=\Big(\prod_{q=1}^{Q}\tfrac{n-k}{n}\Big)\,\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{X^{(j)}\in D\}}$ (5)
is unbiased for any $\xi$ and $k$. It suffices that $D\subset\{\xi\geq I_{\max}\}$, without requiring $\xi(x)\geq I_{\max}\iff x\in D$ [28, 37, 24]. Unbiasedness extends to unnormalised measures $\gamma(\varphi)=\mathbb{E}[\varphi(X_{\tau})\mathbf{1}_{D}(X_{\tau})]$. Randomised cloning and correct handling of ties are necessary to avoid bias.
@@
A law of large numbers holds for AMS estimators. Under mild assumptions,
$\sqrt{n}\Big(\gamma^{(n)}_{1}(\varphi)-\gamma_{1}(\varphi)\Big)\ \Rightarrow\ \mathcal{N}(0,\sigma_{1}^{2}(\varphi)),$
with asymptotic variance characterized via the Fleming–Viot formulation [24, 38]. For $k=1$ and target probability $p$,
@@
A general CLT for $k>1$ remains open, though evidence suggests $n^{-1/2}$ scaling with variance comparable to SMC.
@@
Unbiasedness does not depend on $\xi$, but variance does. Poor or multi-channel choices inflate variance and may yield heavy-tailed errors. In practice, variance is controlled by testing alternative $\xi$ and adjusting $n$ or $k$ [24, 28].
@@
AMS adapts intermediate levels and branching rates on the fly, removing the need for a priori specification as in classical Multilevel Splitting [39] or Sequential Monte Carlo [40, 41]. The algorithm maintains a fixed population size $n$, ensuring robustness, parallel efficiency, and predictable memory use. It provides unbiased estimators for both rare-event probabilities and unnormalised measures $\gamma(\varphi)$, enabling straightforward parallelization across independent runs [24, 28, 37].
@@ 3.1 Antithetic variates: overview and limitations
Antithetic variates [42] reduce variance by pairing negatively correlated samples. In option pricing this corresponds to simulating each path together with its reflection obtained by negating Brownian increments. For monotone payoffs the estimator variance decreases, with a theoretical maximum reduction by a factor of two.
@@ 3.2 Control variates: overview and limitations
$\mathrm{Var}(\hat{\psi}_{\mathrm{CV}})=\frac{1}{n}\mathrm{Var}(Y)(1-\rho^{2}_{Y,W}).$
Variance reduction is therefore effective only when $W$ is strongly correlated with $Y$.For digital or Asian binaries, suitable highly correlated controls are unavailable, and variance reduction is marginal.
@@ 3.3 Multilevel Monte Carlo: overview and limitations
Multilevel Monte Carlo (MLMC) [15, 16] exploits a hierarchy of approximations $X_{0},\dots,X_{L}$ of the same quantity. The telescoping identity
@@
reduces variance by coupling successive levels with shared randomness. The resulting estimator achieves mean-square error $\mathcal{O}(\varepsilon^{2})$ at cost $\mathcal{O}(\varepsilon^{-2})$, compared to $\mathcal{O}(\varepsilon^{-3})$ for standard Monte Carlo [15].
MLMC is effective for standard option pricing but less suited to rare-event estimation. In tail regimes, the variance of inter-level differences decays slowly, limiting efficiency for digital and barrier options. Optimal allocation of samples,
@@
depends on variances $V_{\ell}$ that are themselves costly to estimate and may behave irregularly across levels, especially in rare-event settings. These features complicate implementation and reduce the expected efficiency gains.
@@ 3.4 Importance sampling: overview and limitations
Importance sampling (IS) [43] estimates $\psi=\mathbb{E}[h(X)]$ by sampling from an alternative density $g$ and reweighting:
@@
A common construction is exponential tilting via Girsanov’s theorem. For Brownian-driven models, $g_{\theta}(y)=e^{\theta y-\psi(\theta)}f(y)$ with cumulant generating function $\psi(\theta)=\log\mathbb{E}[e^{\theta Y}]$. The optimal parameter $\theta^{*}$ satisfies $\psi^{\prime}(\theta^{*})=a$, where $a$ is the rare-event threshold.
In rare-event regimes IS becomes unstable. When exercise probabilities are of order $10^{-6}$, the equation $\psi^{\prime}(\theta)=a$ may lack a solution or yield extreme $\theta^{*}$, and evaluation of $e^{\theta Y}$ produces flat likelihood landscapes with sporadic spikes. In such cases Newton–Raphson and related solvers fail to converge, and stochastic optimisers are equally unreliable [44].
@@
Variance explosion: an inappropriate choice of $g(y)$ can inflate the estimator’s variance instead of reducing it [43].
Payoff-specific design: effective importance sampling must be tailored to the payoff. Binary calls, binary puts, and Asian options require distinct tilting schemes, and multi-asset payoffs add combinatorial complexity [13].
AMS can be interpreted as a non-parametric analogue of IS: it requires only an importance function indicating progress toward the rare set, avoiding explicit tilting densities and unstable root-finding, and thus offering broader applicability across option classes.
@@ 4 AMS applications in finance
Having established the theoretical framework, AMS is now applied to binary option pricing under the Black–Scholes and Heston models. The Markov property of both dynamics makes them directly compatible with AMS, which relies on memoryless trajectories. The method is tested on the four binary contracts of Section 2.2, with efficiency gains most evident for deep out-of-the-money options where standard Monte Carlo becomes infeasible.
@@ 4.1 Importance function design
AMS performance depends critically on the importance function $\xi$, which steers trajectories toward the rare-event set. Two constructions are considered:
  • Path-based functions. For European binaries, $\xi$ is the asset price $S_{t}$; for Asian binaries, the arithmetic average up to $t$, $\tfrac{1}{t}\sum_{i=0}^{t}S_{t_{i}}$. For puts, the sign is inverted. In all cases $L_{\max}=K$ ensures $D\subseteq\{\xi>L_{\max}\}$.
@@
    with $d_{2}=\tfrac{\ln(S/K)+(r-\sigma^{2}/2)T}{\sigma\sqrt{T}}$. At each $t$, $S_{t}$ (or the running average for Asians) is inserted as the spot input, regardless of the underlying model. Although exact only for European binaries under Black–Scholes, this construction captures the curvature of the pricing function and improves guidance toward the rare-event region. Here $L_{\max}=0.5$ ensures $D\subseteq\{\xi>L_{\max}\}$.
Lemma 4.1 (Unbiasedness of AMS for digital options).
Let $(S_{t})$ follow either the Black–Scholes dynamics
@@
$\begin{cases}dS_{t}=rS_{t}\,dt+\sqrt{V_{t}}\,S_{t}\,dW_{t}^{(1)},\\[4.0pt] dV_{t}=\kappa(\theta-V_{t})\,dt+\xi\sqrt{V_{t}}\,dW_{t}^{(2)},\end{cases}$
@@
Let $D$ be the rare–event set corresponding to the digital payoff (European or Asian, call or put). For the importance functions $\xi$ introduced in Section 4.1, the sufficient condition
@@
of [28, Theorem 3.2] holds. Then the AMS estimator of the risk–neutral probability $p=\mathbb{Q}(D)$ is
@@
4.2 Parameter setting and option strikes for algorithm performance analysis
For the Black–Scholes model, volatility is fixed at $\sigma=0.2$. For the Heston model, parameters are set to $\rho=-0.5$, $\kappa=2.0$, $\theta=0.04$, and $\psi=0.3$.
Tests include European digital calls and puts with strikes $2.2$ and $0.29$, and Asian digitals. Under Black–Scholes, Asian strikes are $1.7$ (call) and $0.63$ (put); under Heston, $1.6$ and $0.55$. In all cases option values are of order $7.5\times 10^{-6}$, representing rare-event regimes suitable for assessing AMS performance.
4.3 Results and discussion
4.3.1 Impact of the selection parameter K on algorithm performance
The selection parameter $K$ determines the fraction of trajectories discarded at each iteration. Theory shows that asymptotic variance is minimized at $K=1$ with an optimal importance function [24, 38], but this setting is computationally prohibitive.
We examine $K$ values from 5% to 45% of N=50, [^0] N=50{,}000 particles, in 5% increments, for a digital call under Heston and an Asian digital call under Black–Scholes, both using the path-based importance function (Section 4.1).
@@ 4.3.1 Impact of the selection parameter K on algorithm performance
Figure 2: Relationship between $k$ and the normalized variance (horizontal axis) for the simulation results
In the figures, blue markers correspond to the standard digital call option, while red markers represent the Asian digital call option.
Results confirm the trade-off: small $K$ requires more iterations and substantially longer runtime (up to 30 seconds in the Heston case). Estimator quality, however, shows no clear monotonic dependence on $K$, for these options, performance remains stable across the tested range.
4.3.2 Impact of the number of trajectories N on algorithm performance
The particle count $N$ directly affects AMS performance. Larger $N$ reduces estimator variance but increases runtime due to higher simulation and sorting costs. Theoretical analysis shows complexity of order $N\log(p)\log(N)$, accounting for the sorting step and the generation of one new trajectory per iteration [24, 38].
Numerical experiments under both Black–Scholes and Heston models, using the options of Section 4.3.1, confirm this trade-off. All tests use $K=0.45$ and the path-based importance function.
@@ 4.3.2 Impact of the number of trajectories N on algorithm performance
As reported in Table 2, computational cost grows consistently with $N$, in agreement with the predicted $N\log(p)\log(N)$ scaling. Substituting the estimated $p$ and tested $N$ values yields an approximately constant ratio, supporting the theoretical complexity analysis.
@@
4.3.3 Analysis of option pricing results
Standard Monte Carlo is benchmarked against AMS, with multilevel Monte Carlo (MLMC) and antithetic variates as additional baselines. Within AMS, two importance functions are tested. Control variates are excluded due to negligible correlation with the payoff, and importance sampling is omitted since optimal tilting fails to converge in the rare-event regime considered and requires payoff-specific design.
Test cases focus on deep out-of-the-money contracts with exercise probabilities of order $10^{-6}$, where AMS achieves significant gains. For higher probabilities ($p>10^{-3}$), standard Monte Carlo remains competitive and AMS provides only limited advantage.
Performance is evaluated in terms of computational time (horizontal axis) and relative accuracy (vertical axis), defined as $\frac{\sqrt{\mathrm{Var}}}{\mathrm{Mean}},$ with the mean computed over 50 independent runs. Results are reported in Figures 3,4.
The discard fraction is set to $k=0.45$, as smaller values did not yield systematic variance reduction (Section 4.3.1).
Figure 3: Computational time (log scale) as a function of relative accuracy for different simulation methods for the Heston digital call and the Black–Scholes and Heston Asian digital call; numerical values are reported in Tables 5, 7, and 9
Figure 4: Computational time (log scale) as a function of relative accuracy for different simulation methods for the Heston digital put and the Black–Scholes and Heston Asian digital put; numerical values are reported in Tables 6, 8, and 10
The results demonstrate substantial efficiency gains of AMS across all tested settings (Figures 3,4).

### v2

@@ 2.3.1 Theoretical properties of AMS
Well posedness and termination.
Let $X=(X_{t})_{t\geq 0}$ be a Markov process with importance function $\xi$ and rare set $D$. For fixed population size $N$ and killing parameter $K\in\{1,\dots,N-1\}$, AMS is well posed: at each iteration the cutting level $Z$ is an order statistic $Z=S^{(K)}$, and under standard regularity assumptions on $X$ (Feller property), on $\xi$ (continuity), and on the entrance into $D$ (strict entrance condition), the algorithm terminates almost surely after a finite number of iterations [c1].
@@
The AMS estimator
$\hat{p}_{\mathrm{AMS}}=\Big(\tfrac{N-K}{N}\Big)^{Q}\,\frac{1}{N}\sum_{j=1}^{N}\mathbf{1}_{\{X^{j}\in D\}}$ (4)
is unbiased for any admissible choice of $\xi$ and $K$, where $Q$ denotes the (random) number of iterations. It is sufficient that the rare set is contained in a super-level set of the score,
$D\subset\{x:\xi(x)\geq L_{\max}\},$
without requiring the stronger equivalence $\xi(x)\geq L_{\max}\iff x\in D$ [brehier2016unbiasedness, 10.1093/biomet/asy028, c1]. Unbiasedness extends to unnormalised measures
$\gamma(\varphi)\;=\;\mathbb{E}\big[\varphi(X_{\tau})\,\mathbf{1}_{\{X_{\tau}\in D\}}\big],$
for bounded test functions $\varphi$.
A crucial technical point is that randomised cloning and correct handling of times in the scores are essential to preserve unbiasedness. In practice, any deterministic rule for selecting which trajectories to kill or clone breaks exchangeability and introduces bias.
@@
Let $\gamma^{(N)}(\varphi)$ denote the AMS approximation of $\gamma(\varphi)$ based on $N$ replicas. A law of large numbers holds:
$\gamma^{(N)}(\varphi)\ \xrightarrow[N\to\infty]{\mathbb{P}}\ \gamma(\varphi).$
Under additional regularity assumptions, a central limit theorem is available,
$\sqrt{N}\Big(\gamma^{(N)}(\varphi)-\gamma(\varphi)\Big)\ \Rightarrow\ \mathcal{N}\big(0,\sigma^{2}(\varphi)\big),$
where the asymptotic variance $\sigma^{2}(\varphi)$ can be characterised via the associated Fleming–Viot particle system [c1, doi:10.1137/18M1187477]. In particular, for $K=1$ and a rare event with probability $p$,
@@
A general CLT for $K>1$ remains an open problem; however, both theoretical results in simplified settings and numerical evidence indicate the usual $N^{-1/2}$ scaling, with an asymptotic variance comparable to that of Sequential Monte Carlo methods.
@@
While the unbiasedness of $\hat{p}_{\mathrm{AMS}}$ does not depend on the particular choice of the importance function $\xi$, the variance of the estimator is strongly affected by it. In particular, if $\xi$ is poorly adapted to the geometry of the rare event (for example, in analogy with a badly chosen importance sampling density), the resulting estimator may exhibit large variance and therefore a much slower convergence to the true value. In practice, variance is controlled by experimentally comparing alternative definitions of $\xi$ and by suitably adjusting the population size $N$ and the killing parameter $K$ [c1, brehier2016unbiasedness].
@@
AMS adapts intermediate levels and branching rates on the fly, removing the need for a priori specification as in classical Multilevel Splitting [Kahn1951SplittingParticleTransmission] or Sequential Monte Carlo [fba646e2e09d4f7f93e66f71554f16b7, cerou:inria-00071391]. The algorithm maintains a fixed population size $N$, ensuring robustness, parallel efficiency, and predictable memory use. It provides unbiased estimators for both rare event probabilities and unnormalised measures $\gamma(\varphi)$, enabling straightforward parallelization across independent runs [c1, brehier2016unbiasedness, 10.1093/biomet/asy028].
@@ 3.1 Antithetic variates: overview and limitations
Antithetic variates [gentle2009antithetic] reduce variance by pairing negatively correlated samples. In option pricing this corresponds to simulating each path together with its reflection obtained by negating Brownian increments. For monotone payoffs the estimator variance decreases, with a theoretical maximum reduction by a factor of two.
@@ 3.2 Control variates: overview and limitations
$\mathrm{Var}(\hat{\psi}_{\mathrm{CV}})=\frac{1}{n}\mathrm{Var}(Y)\bigl(1-\rho^{2}_{Y,W}\bigr).$
Variance reduction is therefore effective only when $W$ is strongly correlated with $Y$.
In practice, however, there is no unique, model–independent choice of $W$ that works well across products: an effective control variate must be engineered to closely mimic the structure of the specific payoff under the given dynamics. For example, for arithmetic Asian options one can use the optimized lower-bound proxy of [FusaiKyriakou2016]. This makes the technique very powerful when specialized, but at the cost of poor generalizability.
For discontinuous payoffs such as digital or Asian binary options, suitable highly correlated controls are often unavailable in a generic form; without such carefully tailored constructions, the achievable variance reduction remains marginal and the method loses much of its appeal as a general-purpose tool for rare events.
@@ 3.3 Multilevel Monte Carlo: overview and limitations
Multilevel Monte Carlo (MLMC) [m1, m2] exploits a hierarchy of approximations $X_{0},\dots,X_{L}$ of the same quantity. The telescoping identity
@@
reduces variance by coupling successive levels with shared randomness. The resulting estimator achieves mean square error $\mathcal{O}(\varepsilon^{2})$ at cost $\mathcal{O}(\varepsilon^{-2})$, compared to $\mathcal{O}(\varepsilon^{-3})$ for standard Monte Carlo [m1].
MLMC is effective for standard option pricing but less suited to rare event estimation. In tail regimes, the variance of inter-level differences decays slowly, limiting efficiency for digital and barrier options. Optimal allocation of samples,
@@
depends on variances $V_{\ell}$ that are themselves costly to estimate and may behave irregularly across levels, especially in rare event settings. These features complicate implementation and reduce the expected efficiency gains.
@@ 3.4 Importance sampling: overview and limitations
Importance sampling (IS) [tokdar2010importance] estimates $\psi=\mathbb{E}[h(X)]$ by sampling from an alternative density $g$ and reweighting:
@@
A common construction is exponential tilting via Girsanov’s theorem. For Brownian driven models, $g_{\theta}(y)=e^{\theta y-\psi(\theta)}f(y)$ with cumulant generating function $\psi(\theta)=\log\mathbb{E}[e^{\theta Y}]$. The optimal parameter $\theta^{*}$ satisfies $\psi^{\prime}(\theta^{*})=a$, where $a$ is the rare event threshold.
In rare event regimes IS becomes unstable. When exercise probabilities are of order $10^{-6}$, the equation $\psi^{\prime}(\theta)=a$ may lack a solution or yield extreme $\theta^{*}$, and evaluation of $e^{\theta Y}$ produces flat likelihood landscapes with sporadic spikes. In such cases Newton–Raphson and related solvers fail to converge, and stochastic optimisers are equally unreliable [casella2021choice].
@@
Variance explosion: an inappropriate choice of $g(y)$ can inflate the estimator’s variance instead of reducing it [tokdar2010importance].
Payoff specific design: effective importance sampling must be tailored to the payoff. Binary calls, binary puts, and Asian options require distinct tilting schemes, and multi-asset payoffs add combinatorial complexity [imps].
AMS can be interpreted as a non-parametric analogue of IS: it requires only an importance function indicating progress toward the rare set, avoiding explicit tilting densities and unstable root finding, and thus offering broader applicability across option classes.
@@ 4 AMS applications in finance
Having established the theoretical framework, AMS is now applied to binary option pricing under the Black–Scholes and Heston models. The Markov property of both dynamics makes them directly compatible with AMS, which relies on memoryless trajectories. The method is tested on the three binary contracts of Section 2.2, with efficiency gains most evident for deep out-of-the-money options where standard Monte Carlo becomes infeasible.
For the Black–Scholes model, volatility is fixed at $\sigma=0.2$. For the Heston model, parameters are set to $\rho=-0.5$, $\kappa=2.0$, $\theta=0.04$, $\psi=0.3$, $V_{0}=0.04$.
Tests include: a European digital call under Heston dynamics with strike $2.2$; a path-dependent digital up-and-in barrier call under Heston dynamics with barrier level $K_{L}=2.45$; and Asian digital calls under both Black–Scholes (strike $1.7$) and Heston (strike $1.6$). These parameter choices place all contracts in comparable rare-event regimes.
Implementation and hardware.
All numerical experiments were run on a MacBook Pro equipped with an Apple M2 Pro processor and 16 GB of RAM, running macOS. The C++ core of the amsSim library was compiled via Rcpp using the Apple clang compiler with -O3 optimisation flags and no GPU acceleration. All simulations are strictly single-threaded: OpenMP and other forms of multi-threading are disabled, so the reported timings correspond to wall-clock times on a single CPU core.
@@ 4.1 Importance function design
AMS performance depends critically on the importance function $\xi$, which steers trajectories toward the rare event set. Two constructions are considered:
  • Path-based functions. For European binaries, $\xi_{t}=S_{t}$; for Asian binaries, $\xi_{t}$ is the running average; for barrier binaries, $\xi_{t}$ is the running maximum or minimum, depending on the barrier type. For puts, the sign is inverted. In all cases, setting $L_{\max}$ equal to the strike (or to $K_{L}$ for barriers) guarantees $D\subseteq\{\xi>L_{\max}\}$.
@@
    with $d_{2}=\tfrac{\ln(S/K)+(r-\sigma^{2}/2)T}{\sigma\sqrt{T}}$. At each $t$, $S_{t}$ (or the running average for Asians, and the running maximum or minimum for barrier options). is inserted as the spot input, regardless of the underlying model. Although exact only for European binaries under Black–Scholes, this construction captures the curvature of the pricing function and improves guidance toward the rare event region. Here $L_{\max}=0.5$ ensures $D\subseteq\{\xi>L_{\max}\}$.
Remark 4.1 (Unbiasedness of AMS for digital options).
Let $(S_{t})$ follow either the Black-Scholes dynamics
@@
$\begin{cases}dS_{t}=rS_{t}\,dt+\sqrt{V_{t}}\,S_{t}\,dW_{t}^{(1)},\\[4.0pt] dV_{t}=\kappa(\theta-V_{t})\,dt+\psi\sqrt{V_{t}}\,dW_{t}^{(2)},\end{cases}$
@@
Let $D$ be the rare–event set corresponding to the digital payoff (European, Asian or barrier call or put). For the importance functions $\xi$ introduced in Section 4.1, the sufficient condition
@@
of [brehier2016unbiasedness, Theorem 3.2] holds. Then the AMS estimator of the risk–neutral probability $p=\mathbb{Q}(D)$ is
@@
4.1.1 Heuristics and diagnostics for importance function design
From a practical standpoint, the importance function $\xi$ should be chosen so that it is monotonically aligned with the event of interest. Writing the rare event as $D=\{x:G(x)>0\}$ for a suitable functional $G$ of the path, a natural heuristic is to take $\xi(x)$ as a smooth surrogate of $G(x)$: trajectories closer to $D$ (in the sense of $G(x)$ being larger) should also have larger $\xi(x)$. In our constructions this is enforced explicitly: for digital payoffs we use functionals that increase as the underlying approaches the exercise region (terminal value, running average, running extremum).
From a theoretical perspective, a first requirement is that $\xi$ preserves the support of the event, in the sense that $x\in D\;\;\Longrightarrow\;\;\xi(x)\geq L_{\max},$ so that the sufficient condition of [brehier2016unbiasedness, Theorem 3.2] is satisfied. If this condition were violated, AMS could in principle miss part of $D$ and introduce bias. Even when unbiasedness holds, a poorly informative choice of $\xi$ (for instance, a score that is nearly uncorrelated with $\mathbf{1}_{D}$ or that does not increase as trajectories move towards $D$) inflates the variance and can lead to worse efficiency than crude Monte Carlo.
4.2 Impact of the selection parameter $K$ on algorithm performance
The selection parameter can be expressed either as the absolute number of discarded trajectories $K$, or equivalently as the discard fraction $k=K/N$, where $N$ is the population size. At each iteration exactly $K=kN$ particles (i.e., a fraction $k$ of the population) are killed and replaced by clones. Theoretical results show that, for an optimal importance function, the asymptotic variance is minimised in the limit $K=1$ [c1, doi:10.1137/18M1187477], but this regime is computationally impractical in realistic settings.
In our experiments we fix N=50, [^0] N=50{,}000 particles and vary the discard fraction $k$ from 5% to 45% in steps of 5%, so that $K=kN$ trajectories are removed at each iteration. We investigate this dependence for a digital call under Heston dynamics and for an Asian digital call under Black–Scholes, both using the path-based importance function described in Section 4.1.
@@ 4.2 Impact of the selection parameter $K$ on algorithm performance
Figure 2: Normalized variance (vertical axis) as a function of the discard fraction k (horizontal axis).
In the figure, gray markers correspond to the standard digital call option, while red markers represent the Asian digital call option.
Results confirm the trade-off: small $K$ requires more iterations and substantially longer runtime (up to 30 seconds in the Heston case). Estimator quality, however, shows no clear monotonic dependence on K; for these options, performance remains stable across the tested range.
4.3 Impact of the number of trajectories N on algorithm performance
The particle count $N$ directly affects AMS performance. Larger $N$ reduces estimator variance but increases runtime due to higher simulation and sorting costs. Theoretical analysis shows complexity of order $N(-\log(p))\log(N)$, accounting for the sorting step and the generation of one new trajectory per iteration [c1, doi:10.1137/18M1187477].
Numerical experiments under both Black–Scholes and Heston models, using the options of Section 4.2, confirm this trade-off. All tests use $K=0.45$ and the path-based importance function.
@@ 4.3 Impact of the number of trajectories N on algorithm performance
As reported in Table 2, the computational cost increases steadily with $N$, as expected. To compare with the predicted $N(-\log p)\log N$ complexity, we consider the ratio between the observed execution time and $N\log N(-\log p)$.
In our experimental settings for the Heston digital option this ratio is approximately $1.1\times 10^{-5}(-\log p)$ across all tested values of $N$, while for the Black–Scholes Asian digital option it is about $7.5\times 10^{-6}(-\log p)$, indicating an essentially constant prefactor and thus supporting the theoretical complexity analysis.
@@
4.4 Analysis of option pricing results
Standard Monte Carlo serves as the primary benchmark for all contracts, together with antithetic variates, multilevel Monte Carlo (MLMC), and adaptive multilevel splitting (AMS) using two different importance functions. These constitute the core comparison set across all experiments. In addition, for specific payoffs further specialised methods are included when appropriate, enabling a contract-by-contract performance assessment.
Test cases focus on deep out-of-the-money contracts, with exercise probabilities down to order $10^{-6}$, where AMS is expected to deliver its largest efficiency gains. In our experiments, substantial speed-ups over crude Monte Carlo only start to appear once the probability drops below roughly $10^{-4}$; for less extreme strikes, standard Monte Carlo remains competitive and AMS offers at best modest improvements. These moderately rare events are therefore not the primary target of the present study.
Performance is evaluated in terms of computational time (horizontal axis) and relative accuracy (vertical axis), defined as
$\frac{\sqrt{\mathrm{Var}}}{\mathrm{Mean}},$
with the mean and variance estimated over $50$ independent runs. Figures 3 and 4 summarise these results: each curve represents one simulation method for a given contract, and each point on a curve corresponds to a different computational budget. Times are shown on a logarithmic scale, so vertical separations between curves can be interpreted directly as orders of magnitude in speed-up. The underlying numerical values used to generate these plots are reported in Tables 69.
The discard fraction is set to $k=0.45$ (so that $K=0.45N$ trajectories are resampled at each iteration), as smaller values did not yield systematic variance reduction for the same total computational budget (Section 4.2).
We denote by MCA Monte Carlo with antithetic variates, and by MCV Monte Carlo with the optimised lower-bound control variate of Fusai and Kyriakou [FusaiKyriakou2016].
Figure 3: Computational time (log scale) as a function of relative accuracy for different simulation methods for the Heston digital call and digital up-and-in barrier call; numerical values are reported in Tables 6, 7
Figure 4: Computational time (log scale) as a function of relative accuracy for different simulation methods for the Black-Scholes and Heston digital Asian call; numerical values are reported in Tables 8, 9
Figures 3 and 4 show that, across all contracts considered, the AMS curves lie strictly below the corresponding Monte Carlo and MLMC curves over the range of relative accuracies of interest. In other words, for any target relative error in the plots, AMS attains that accuracy at a lower computational cost. This visual separation is most pronounced in the rare-event regime, where standard Monte Carlo must simulate a very large number of paths before observing enough exercise events.
In Figure 4, an additional benchmark is included: the control variate method derived from the optimised lower bound introduced by [FusaiKyriakou2016]. This provides a further point of comparison, confirming that even when specialised variance–reduction techniques tailored to the payoff structure are incorporated, AMS remains the most computationally efficient method across the accuracy levels examined.
Computational time reduction. For the standard European digital option, AMS achieves speedups exceeding $100\times$ relative to standard Monte Carlo, peaking above $200\times$ around the $5\%$ relative accuracy level. The gap between the AMS and MLMC curves remains close to two orders of magnitude throughout this regime. A similarly pronounced effect is observed for the digital up-and-in barrier option under Heston: AMS retains speedups consistently above $100\times$ even when compared with the best-performing baseline, namely MLMC.